Circular No. 7123

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 C1 (SOHO) D. A. Biesecker, NASA Goddard Space Flight Center, reports obser-

vations of a probable Kreutz sungrazing comet in SOHO/LASCO C3 and C2 data on Feb. 6 and 7. The comet was a relatively bright one, with a tail over 0^{o}.25 long, and it was discovered by D. Lewis, University of Birmingham, in C3 data at a projected heliocentric distance of ∼ 25 R_{⊙}. The last observation was in C2 data some 28 hr later at ∼ 3.9 R_{⊙}.

1999 UT  α_{2000}  δ_{2000}
Feb. 6.459  21^{h}49.^{m}2  −17^{o}59′

Astrometric measurements, made by Biesecker and S. Gregory and reduced by G. V. Williams, are given on MPEC 1999-E18, together with orbital elements by B. G. Marsden.

XTE J1550–564 R. Remillard, E. Morgan, and A. Levine, Massachusetts Institute of

Technology; J. McClintock and G. Sobczak, Center for Astrophysics; C. Bailyn and R. Jain, Yale University; and J. Orosz, Pennsylvania State University, write: “There is continued evolution of the high-frequency quasiperiodic oscillations (QPO) in the x-ray nova and blackhole candidate XTE J1550–564 (see IAUC 7121). An RXTE PCA observation on Mar. 5.52 UT reveals a QPO at 283 ± 5 Hz with a FWHM of 39 ± 9 Hz, and an rms amplitude of 4.0 ± 0.7 percent at 6–30 keV, similar to that reported for Mar. 4.80. However, subsequent PCA observations show changes as follows: 235 ± 6 Hz (FWHM 30 ± 11; amplitude 2.8 ± 0.8 percent at 6–30 keV) on Mar. 7.50; 182.0 ± 2.5 Hz (FWHM 12 ± 4; amplitude 0.91 ± 0.23 percent at 2–30 keV) on Mar. 8.37. Thus the frequency has returned to the value at which the QPO was originally discovered (IAUC 7025). The steep increase in amplitude with photon energy (IAUC 7121) has lessened on Mar. 8, as the QPO is weakly visible in the band 2–6 keV, while the profile fits have poor quality when the energy range is restricted to 6–30 keV.”

COMET C/1998 P1 (WILLIAMS) Visual m_{1} estimates: Feb. 6.61 UT, 9.2 (S. Yoshida, Ibaraki, Japan,

0.25-m reflector); 18.95, 9.5 (J. Carvajal, Madrid, Spain, 0.21-m reflector); 25.15, 10.7 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); Mar. 6.78, 10.9 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); 8.82, 10.9 (M. Meyer, Frauenstein, Germany, 0.25-m reflector).

1999 March 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green